# Action Reasoning Models that can Reason in Space

### v2

@@ Abstract
Reasoning is central to purposeful action, yet most robotic foundation models map perception and instructions directly to control, which limits adaptability, generalization, and semantic grounding. We introduce Action Reasoning Models (ARMs), a class of robotic foundation models that integrates perception, planning, and control through a structured three-stage pipeline. Our model, MolmoAct, encodes observations and instructions into depth-aware perception tokens, generates mid-level spatial plans as editable trajectory traces, and predicts precise low-level actions, enabling explainable and steerable behavior. MolmoAct-7B-D achieves strong performance across simulation and real-world settings: $70.5\%$ zero-shot accuracy on SimplerEnv Visual Matching tasks, surpassing closed-source $\pi_{0}$ and GR00T N1; $86.6\%$ average success on LIBERO, including a $6.3\%$ gain over ThinkAct on long-horizon tasks; and in real-world fine-tuning, $10\%$ (single-arm) and $22.7\%$ (bimanual) task progression over $\pi_{0}$-FAST. It also outperforms baselines by $23.3\%$ on out-of-distribution generalization and achieves top human-preference scores for open-ended instruction following and trajectory steering. Furthermore, we release, for the first time, the MolmoAct Dataset —a mid-training robot dataset comprising over 10,000 high-quality robot trajectories across diverse scenarios and tasks. Training with this dataset yields an average 5.5% improvement in general performance over the base model. We release all model weights, training code, MolmoAct Dataset and our action reasoning dataset, establishing MolmoAct as both a state-of-the-art robotics foundation model and an open blueprint for building ARMs that transform perception into purposeful action through grounded reasoning.
@@ 1 Introduction
MolmoAct’s structured design delivers both strong performance and high explainability. On standard benchmarks such as LIBERO and SimplerEnv (Google Robot), MolmoAct consistently outperforms competitive baselines including GR00T N1 (NVIDIA et al., 2025), $\pi_{0}$ and $\pi_{0}$-FAST (Black et al., ), RT-1 (Brohan et al., 2022), and TraceVLA (Zheng et al., 2024). In arena-style human evaluations for open-ended language instruction following, MolmoAct is preferred over baselines, achieving significantly higher Elo ratings. The model adapts to novel tasks more effectively through lightweight fine-tuning, surpassing other strong baselines in efficiency. Moreover, it generalizes well to diverse environments and task perturbations in both simulation and real-world settings. Its visual reasoning traces offer an explainable view into the model’s decision-making, while also enabling direct action steering by editing trajectory lines—an approach we find more reliable than language commands, which can suffer from ambiguity.
@@ 2.2 Vision-Language-Action Model
Prior work represents the 256 discretized action bins with 256 distinct language tokens taken from the tail of the vocabulary. Continuous action bins, however, possess ordinal structure and local correlation, whereas arbitrary language tokens are effectively unrelated. This mismatch yields a poor initialization for learning an action codebook. We adopt a simple alternative that better reflects the geometry of the action space. We first identify the final 256 tokens in the Qwen2 tokenizer and, for each, use its underlying byte-level BPE symbol. We then assign them monotonically to the 256 bins so that adjacent bins map to adjacent symbols, and this becomes our action token vocabulary $V_{\mathrm{action}}$. We notice that these BPE symbols have a better initialization for action token embeddings by sharing similar characters between adjacent action bins. This similarity-preserving initialization offers a smoother starting point for optimization and, in practice, substantially reduces training time. In contrast to GR00T N1’s (NVIDIA et al., 2025) training time of 50,000 GPU hours, MolmoAct achieves pre-training with only 9,216 GPU hours: over a 5x reduction.
@@ 2.3 Action Reasoning Model
$V_{\mathrm{depth}}=\{\langle\mathrm{DEPTH\_START}\rangle,\langle\mathrm{DEPTH\_END}\rangle\}\;\cup\;\{\langle\mathrm{DEPTH}\_k\rangle\}_{k=1}^{N}$ (1)
@@
$\mathbf{d}=\big{(}\langle\mathrm{DEPTH\_START}\rangle\;\langle\mathrm{DEPTH}\_{z^{\mathrm{depth}}_{1}}\rangle\;\ldots\;\langle\mathrm{DEPTH}\_{z^{\mathrm{depth}}_{M}}\rangle\;\langle\mathrm{DEPTH\_END}\rangle\big{)}\in V_{\mathrm{depth}}.$ (2)
with $M=100$, and each $z^{\mathrm{depth}}_{i}\in\{1,\ldots,N\}$ indexing a code in a VQVAE (Van Den Oord et al., 2017) codebook $\mathcal{C}=\{c_{1},\ldots,c_{N}\}$. The codebook is produced by a pre-trained specialist depth estimator (trained on depth maps from Depth Anything V2), which quantizes the dense depth map into a fixed-length sequence of $M$ indices. There is a deterministic one-to-one correspondence between each codebook index and a depth token in $V_{\mathrm{depth}}$ (i.e., index $k$ maps to $\langle\mathrm{DEPTH}\_k\rangle$), so $\mathbf{d}$ is a discrete, explainable summary of the depth map of the scene. We employ a specialist-to-generalist distillation strategy to ground MolmoAct’s depth perception tokens prediction: the specialist produces $\mathbf{d}$ as the ground-truth depth string, and the VLA is trained to predict this string autoregressively from the original RGB observation, thereby internalizing depth in a form that can condition downstream trajectory and action generation.
@@
Given a certain image observation, we define the end-effector visual reasoning trace on that image as a polyline with $L$ points, $1\!\leq\!L\!\leq\!5$ (i.e., 0 to 4 line segments),
@@
$p(\mathbf{d},\boldsymbol{\tau},\mathbf{a}\,\big{\lvert}\,I,T)=\>\prod_{i=1}^{M+2}p\bigl{(}d_{i}\mid I,T,\mathbf{d}_{<i}\bigr{)}\;\times\;\prod_{j=1}^{L}p\bigl{(}\tau_{j}\mid I,T,\mathbf{d},\boldsymbol{\tau}_{<j}\bigr{)}\;\times\;\prod_{k=1}^{D}p\bigl{(}a_{k}\mid I,T,\mathbf{d},\boldsymbol{\tau},\mathbf{a}_{<k}\bigr{)}.$ (4)
@@ 2.4 Action Steerability via Visual Reasoning Trace
We define steerability as the ability to guide a policy at test time to perform different behaviors using user-provided instructions. Most prior VLA systems rely exclusively on language for steering. However, language-only steering faces three practical challenges: (i) it requires large and diverse corpora of high-quality language–action pairs to learn a reliable grounding between words and control, (ii) natural language is inherently ambiguous about magnitudes, scales, and endpoints, and (iii) post-trained models often exhibit narrow prompting habits, making them brittle to out-of-distribution phrasing. For manipulation, these issues translate into imprecise or inconsistent control. We therefore seek a steering modality that is both precise and scalable. Rather than relying on ambiguous language prompts, we allow the user to draw a visual reasoning trace $\boldsymbol{\tau}$ directly on the camera image to indicate the desired end-effector path. A trace $\boldsymbol{\tau}=(p_{1},\dots,p_{L})$, $1\leq L\leq 5$ (i.e., 0 to 4 line segments), is overlaid onto the RGB image $I$ to form an augmented observation $I^{+}=I\oplus\boldsymbol{\tau}$. visual reasoning traces are unambiguous, easily edited, and generalize across tasks without large text–action corpora or brittle language patterns.
@@ 3.1 Action Reasoning Data
To generate Depth Perception Tokens for each frame of a demonstration, we first train a VQVAE on 10 million depth maps of tabletop manipulation images collected from the RT-1, BridgeData V2, and BC-Z datasets. We use DepthAnything-v2 to obtain a depth map for each observation RGB image. The VQVAE is trained with a standard reconstruction objective to minimize reconstruction loss between input RGB images and their corresponding depth maps for 20 epochs. Once the VQVAE has been trained, we encode each observation image with the VQVAE to get their latent embeddings. We then represent the latent embedding with a learned codebook with 128 dimension based on a one-to-one index to depth token mapping. Note that all images are resized to $\times320$ px during training and inference to enforce the representation of 100 tokens per image. This allows us to express the depth map of each observation image as a tokenized string of 100 tokens, which we use for ground truth labeling for our depth perception token.
@@ 3.2 MolmoAct Dataset
GR00T N1 (fine-tuned) (NVIDIA et al., 2025) 0.7% 1.9% 2.9% 1.8%
@@ 4.2 Mid-training
Following the initial pre-training stage, we conduct a second stage of mid-training using high-quality action reasoning data closely aligned with our target domain of household manipulation. Specifically, we formulate the MolmoAct Dataset into 1M action reasoning data samples and an additional 1M trajectory-conditioned action data samples, which we find beneficial for improving overall performance and action steering. To train on the MolmoAct Dataset, we convert each sample—consisting of two side-mounted camera views and one wrist camera view, all sharing the same instruction and ground-truth action—into two paired-view training examples by pairing each side view with the wrist view. For action reasoning data preparation (subsection 3.1), depth perception tokens and visual reasoning traces are generated only from the side views, while the wrist view is used solely for providing additional information. We train the model on this modified dataset for 50k gradient steps with a batch size of 128, using 128 H100 GPUs over around 2,304 GPU hours. Additional hyperparameters and optimization details are provided in Appendix B.
Table 2: LIBERO benchmark success rates across four task categories (Spatial, Object, Goal, and Long-horizon) along with the average performance. MolmoAct achieves the highest overall average success rate of 86.6%, outperforming all baselines, with strong performance across all categories, particularly in long-horizon tasks.
@@ 4.3 Post-training
After mid-training, we conduct the final post-training stage to rapidly adapt the model to new tasks and embodiments. For new tasks, we collect a small set of 30 to 50 tele-operated demonstrations per task, then generate perception tokens and visual reasoning trace for each timestep. These demonstrations are converted into the action reasoning data and trajectory-conditioned action data, following the same process as mid-training data curation, with one key difference: we apply action chunking (Zhao et al., 2023) during post-training, formatting action predictions in fixed-length chunks ($N=8$). For each action chunk in all chunks, we tokenize each of them in the same way as we do for single action. After grouping them to a list, we train the model autoregressively for predicting all action chunks. We adapt MolmoAct to post-train on single or multi-tasks via parameter-efficient LoRA fine-tuning. In all evaluations, we fix the LoRA rank at 32 and alpha at 16 to preserve the model’s pre-trained capabilities. For simulation benchmarks (e.g., LIBERO), we use a batch size of 128, and for real-world tasks we use 64. The number of gradient steps we train varies by task. Our post-training data generally consists of a front- or side-view image paired with a wrist-view image, although some setups provide multiple wrist views (e.g., bimanual scenarios). In all cases, we apply the same LoRA and training configuration described above. Additional details are available in Appendix B.
@@ 5 Experimental Evaluation
Our experimental evaluation comprises a broad suite of studies that rigorously benchmark MolmoAct against strong baselines. We assess MolmoAct with MolmoAct-7B-D version in (i) its pre-training, “out-of-the-box” capabilities, (ii) its post-training adaptability across varied tasks, domains, and robotic embodiments, and (iii) its additional feature of being an interactive and steerable action reasoning model. By testing the model on a comprehensive range of scenarios both in simulation and real-world, we aim to answer the following research questions:
@@ 5.1 MolmoAct After Pre-training
Evaluation Setup and Baselines. We first evaluated MolmoAct’s zero-shot capability—its ability immediately after pre-training and before any task-specific fine-tuning. Unlike $\pi_{0}$ (Black et al., ), GR00T N1 (NVIDIA et al., 2025), and other proprietary VLA models that rely on large-scale private robot datasets and the full OXE dataset for pre-training, MolmoAct was trained exclusively on curated action reasoning data filtered from a subset of OXE (specifically BC-Z (Jang et al., 2022), BridgeData V2 (Walke et al., 2023), and RT-1 (Brohan et al., 2022)), combined with multimodal web data and auxiliary robot data, as detailed in subsection 4.1. This amounts to approximately 26.3M samples—an order of magnitude smaller than $\pi_{0}$, which uses at least 903M for pre-training. To evaluate MolmoAct’s out-of-the-box generalization, we used the SimplerEnv benchmark, which features both visual-matching and variant-aggregation tasks across WidowX and Google Robot platforms (details in subsection D.1). As MolmoAct ’s pre-training distribution is most aligned with the Google Robot visual-matching tasks, we focused our evaluation on this suite to best isolate in-distribution performance and capabilities of pre-training.
We compared MolmoAct-7B-D-Pretrain against a set of generalist policies, including TraceVLA (Zheng et al., 2024), RT-1X (Brohan et al., 2022), OpenVLA (Kim et al., 2024), RoboVLM (Liu et al., 2025), Emma-x (Sun et al., 2024), $\pi_{0}$ and $\pi_{0}$-FAST (Black et al., ), Octo (Team et al., 2024b), Magma (Yang et al., 2025b), HPT (Wang et al., 2024a), SpatialVLA (Qu et al., 2025) and GR00T N1 (NVIDIA et al., 2025). Most baselines were evaluated in the zero-shot setting, with a subset also tested after fine-tuning. We additionally fine-tuned MolmoAct-7B-D-Pretrain on the RT-1 subset of OXE to assess its capacity when given more pre-training data.
Evaluation Results. MolmoAct-7B-D-Pretrain achieved strong zero-shot performance on the SimplerEnv visual-matching suite, reaching 70.5% success rate and outperforming baselines such as GR00T N1, $\pi_{0}$, $\pi_{0}$-FAST, and Magma. With fine-tuning on the same RT-1 subset of OXE, MolmoAct-7B-D improved to 71.6%, exceeding Magma by 3.2% as shown in Table 1. These results indicate that MolmoAct is both an effective zero-shot generalist and a strong initialization for fine-tuned deployment.
@@ 5.2 Fast Adaptation of MolmoAct in Post-training
Evaluation Setups and Baselines. We evaluate MolmoAct in both simulation and real-world settings to assess its fast adaptation after post-training. In simulation, we evaluate on the LIBERO simulation benchmark (Liu et al. (2023a)), which consists of a Franka Emika Panda arm in simulation with demonstrations containing front and wrist view camera images ($\times256$ px), language instructions, and delta end-effector pose actions. We follow prior works (Kim et al. (2024)) and evaluate on four task suites – LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long – each with 500 demonstrations across 10 tasks. Following (Kim et al. (2024)), we trained on a modified dataset that filtered out no-op actions and unsuccessful demonstrations. Moreover, we set action chunk size to $K=8$ for evaluation on each task suite and execute full chunks before redoing action reasoning. We fine-tune MolmoAct-7B-D using Low-Rank Adaptation (LoRA) and compared to state-of-the-art generalist autoregressive policies, such as TraceVLA (Zheng et al., 2024), OpenVLA (Kim et al., 2024), SpatialVLA (Qu et al., 2025), $\pi_{0}$-FAST (Black et al., ), CoT-VLA (Zhao et al., 2025), WorldVLA (Cen et al., 2025), ThinkAct (Huang et al., 2025), and NORA-AC (Hung et al., 2025).

### v4

@@ Abstract
Reasoning is central to purposeful action, yet most robotic foundation models map perception and instructions directly to control, which limits adaptability, generalization, and semantic grounding. We introduce Action Reasoning Models (ARMs), a class of robotic foundation models that integrates perception, planning, and control through a structured three-stage pipeline. Our model, MolmoAct, encodes observations and instructions into depth-aware perception tokens, generates mid-level spatial plans as editable trajectory traces, and predicts precise low-level actions, enabling explainable and steerable behavior. MolmoAct-7B-D achieves strong performance across simulation and real-world settings: $70.5\%$ zero-shot accuracy on SimplerEnv Visual Matching tasks, surpassing closed-source $\pi_{0}$ and GR00T N1.5; $86.6\%$ average success on LIBERO, including a $6.3\%$ gain over ThinkAct on long-horizon tasks; and in real-world fine-tuning, $10\%$ (single-arm) and $22.7\%$ (bimanual) task progression over $\pi_{0}$-FAST. It also outperforms baselines by $23.3\%$ on out-of-distribution generalization and achieves top human-preference scores for open-ended instruction following and trajectory steering. Furthermore, we release, for the first time, the MolmoAct Dataset —a mid-training robot dataset comprising over 10,000 high-quality robot trajectories across diverse scenarios and tasks. Training with this dataset yields an average 5.5% improvement in general performance over the base model. We release all model weights, training code, MolmoAct Dataset and our action reasoning dataset, establishing MolmoAct as both a state-of-the-art robotics foundation model and an open blueprint for building ARMs that transform perception into purposeful action through grounded reasoning.
[…]
@@ 1 Introduction
MolmoAct’s structured design delivers both strong performance and high explainability. On standard benchmarks such as LIBERO and SimplerEnv (Google Robot), MolmoAct consistently outperforms competitive baselines including GR00T N1.5 (NVIDIA et al., 2025), $\pi_{0}$ and $\pi_{0}$-FAST (Black et al., ), RT-1 (Brohan et al., 2022), and TraceVLA (Zheng et al., 2024). In arena-style human evaluations for open-ended language instruction following, MolmoAct is preferred over baselines, achieving significantly higher Elo ratings. The model adapts to novel tasks more effectively through lightweight fine-tuning, surpassing other strong baselines in efficiency. Moreover, it generalizes well to diverse environments and task perturbations in both simulation and real-world settings. Its visual reasoning traces offer an explainable view into the model’s decision-making, while also enabling direct action steering by editing trajectory lines—an approach we find more reliable than language commands, which can suffer from ambiguity.
@@ 2.2 Vision-Language-Action Model
Prior work represents the 256 discretized action bins with 256 distinct language tokens taken from the tail of the vocabulary. Continuous action bins, however, possess ordinal structure and local correlation, whereas arbitrary language tokens are effectively unrelated. This mismatch yields a poor initialization for learning an action codebook. We adopt a simple alternative that better reflects the geometry of the action space. We first identify the final 256 tokens in the Qwen2 tokenizer and, for each, use its underlying byte-level BPE symbol. We then assign them monotonically to the 256 bins so that adjacent bins map to adjacent symbols, and this becomes our action token vocabulary $V_{\mathrm{action}}$. We notice that these BPE symbols have a better initialization for action token embeddings by sharing similar characters between adjacent action bins. This similarity-preserving initialization offers a smoother starting point for optimization and, in practice, substantially reduces training time. In contrast to GR00T N1.5’s (NVIDIA et al., 2025) training time of 50,000 GPU hours, MolmoAct achieves pre-training with only 9,216 GPU hours: over a 5x reduction.
@@ 2.3 Action Reasoning Model
[…]
@@
[…]
@@
Given a certain image observation, we define the end-effector visual reasoning trace on that image as a polyline with $L$ points, $1\!\le\!L\!\le\!5$ (i.e., 0 to 4 line segments),
@@
[…]
@@ 2.4 Action Steerability via Visual Reasoning Trace
We define steerability as the ability to guide a policy at test time to perform different behaviors using user-provided instructions. Most prior VLA systems rely exclusively on language for steering. However, language-only steering faces three practical challenges: (i) it requires large and diverse corpora of high-quality language–action pairs to learn a reliable grounding between words and control, (ii) natural language is inherently ambiguous about magnitudes, scales, and endpoints, and (iii) post-trained models often exhibit narrow prompting habits, making them brittle to out-of-distribution phrasing. For manipulation, these issues translate into imprecise or inconsistent control. We therefore seek a steering modality that is both precise and scalable. Rather than relying on ambiguous language prompts, we allow the user to draw a visual reasoning trace $\boldsymbol{\tau}$ directly on the camera image to indicate the desired end-effector path. A trace $\boldsymbol{\tau}=(p_{1},\dots,p_{L})$, $1\le L\le 5$ (i.e., 0 to 4 line segments), is overlaid onto the RGB image $I$ to form an augmented observation $I^{+}=I\oplus\boldsymbol{\tau}$. visual reasoning traces are unambiguous, easily edited, and generalize across tasks without large text–action corpora or brittle language patterns.
@@ 3.1 Action Reasoning Data
To generate Depth Perception Tokens for each frame of a demonstration, we first train a VQVAE on 10 million depth maps of tabletop manipulation images collected from the RT-1, BridgeData V2, and BC-Z datasets. We use DepthAnything-v2 to obtain a depth map for each observation RGB image. The VQVAE is trained with a standard reconstruction objective to minimize reconstruction loss between input RGB images and their corresponding depth maps for 20 epochs. Once the VQVAE has been trained, we encode each observation image with the VQVAE to get their latent embeddings. We then represent the latent embedding with a learned codebook with 128 dimension based on a one-to-one index to depth token mapping. Note that all images are resized to 320320 px during training and inference to enforce the representation of 100 tokens per image. This allows us to express the depth map of each observation image as a tokenized string of 100 tokens, which we use for ground truth labeling for our depth perception token.
@@ 3.2 MolmoAct Dataset
GR00T N1.5 (fine-tuned) NVIDIA et al. (2025) 69.3% 68.7% 35.8% 52.4% 46.7% 62.9% 17.5% 43.7%
@@ 4.2 Mid-training
Following the initial pre-training stage, we conduct a second stage of mid-training using high-quality action reasoning data closely aligned with our target domain of household manipulation. Specifically, we formulate the MolmoAct Dataset into 1M action reasoning data samples and an additional 1M trajectory-conditioned action data samples, which we find beneficial for improving overall performance and action steering. To train on the MolmoAct Dataset, we convert each sample—consisting of two side‐mounted camera views and one wrist camera view, all sharing the same instruction and ground‐truth action—into two paired‐view training examples by pairing each side view with the wrist view. For action reasoning data preparation (subsection 3.1), depth perception tokens and visual reasoning traces are generated only from the side views, while the wrist view is used solely for providing additional information. We train the model on this modified dataset for 50k gradient steps with a batch size of 128, using 128 H100 GPUs over around 2,304 GPU hours. Additional hyperparameters and optimization details are provided in Appendix B.
Table 2: LIBERO benchmark success rates across four task categories (Spatial, Object, Goal, and Long-horizon) along with the average performance. MolmoAct achieves the highest overall average success rate of 86.6%, outperforming all autoregressive baselines, with strong performance across all categories, particularly in long-horizon tasks.
@@ 4.3 Post-training
After mid‐training, we conduct the final post‐training stage to rapidly adapt the model to new tasks and embodiments. For new tasks, we collect a small set of 30 to 50 tele‐operated demonstrations per task, then generate perception tokens and visual reasoning trace for each timestep. These demonstrations are converted into the action reasoning data and trajectory‐conditioned action data, following the same process as mid‐training data curation, with one key difference: we apply action chunking (Zhao et al., 2023) during post‐training, formatting action predictions in fixed‐length chunks ($N=8$). For each action chunk in all chunks, we tokenize each of them in the same way as we do for single action. After grouping them to a list, we train the model autoregressively for predicting all action chunks. We adapt MolmoAct to post-train on single or multi-tasks via parameter-efficient LoRA fine-tuning. In all evaluations, we fix the LoRA rank at 32 and alpha at 16 to preserve the model’s pre-trained capabilities. For simulation benchmarks (e.g., LIBERO), we use a batch size of 128, and for real-world tasks we use 64. The number of gradient steps we train varies by task. Our post-training data generally consists of a front- or side-view image paired with a wrist-view image, although some setups provide multiple wrist views (e.g., bimanual scenarios). In all cases, we apply the same LoRA and training configuration described above. Additional details are available in Appendix B.
@@ 5 Experimental Evaluation
Our experimental evaluation comprises a broad suite of studies that rigorously benchmark MolmoAct against strong baselines. We assess MolmoAct with MolmoAct-7B-D version in (i) its pre‑training, “out‑of‑the‑box” capabilities, (ii) its post‑training adaptability across varied tasks, domains, and robotic embodiments, and (iii) its additional feature of being an interactive and steerable action reasoning model. By testing the model on a comprehensive range of scenarios both in simulation and real-world, we aim to answer the following research questions:
@@ 5.1 MolmoAct After Pre-training
Evaluation Setup and Baselines. We first evaluated MolmoAct’s zero-shot capability—its ability immediately after pre-training and before any task-specific fine-tuning. Unlike $\pi_{0}$ (Black et al., ), GR00T N1.5 (NVIDIA et al., 2025), and other proprietary VLA models that rely on large-scale private robot datasets and the full OXE dataset for pre-training, MolmoAct was trained exclusively on curated action reasoning data filtered from a subset of OXE (specifically BC-Z (Jang et al., 2022), BridgeData V2 (Walke et al., 2023), and RT-1 (Brohan et al., 2022)), combined with multimodal web data and auxiliary robot data, as detailed in subsection 4.1. This amounts to approximately 26.3M samples—an order of magnitude smaller than $\pi_{0}$, which uses at least 903M for pre-training. To evaluate MolmoAct’s out-of-the-box generalization, we used the SimplerEnv benchmark, which features both visual-matching and variant-aggregation tasks across WidowX and Google Robot platforms (details in subsection D.1). As MolmoAct ’s pre-training distribution is most aligned with the Google Robot visual-matching tasks, we focused our evaluation on this suite to best isolate in-distribution performance and capabilities of pre-training.
We compared MolmoAct-7B-D-Pretrain against a set of generalist policies, including TraceVLA (Zheng et al., 2024), RT-1X (Brohan et al., 2022), OpenVLA (Kim et al., 2024), RoboVLM (Liu et al., 2025), Emma-x (Sun et al., 2024), $\pi_{0}$ and $\pi_{0}$-FAST (Black et al., ), Octo (Team et al., 2024b), Magma (Yang et al., 2025b), HPT (Wang et al., 2024a), SpatialVLA (Qu et al., 2025) and GR00T N1.5 (NVIDIA et al., 2025). Most baselines were evaluated in the zero-shot setting, with a subset also tested after fine-tuning. We additionally fine-tuned MolmoAct-7B-D-Pretrain on the RT-1 subset of OXE to assess its capacity when given more pre-training data.
Evaluation Results. MolmoAct-7B-D-Pretrain achieved strong zero-shot performance on the SimplerEnv visual-matching suite, reaching 70.5% success rate and outperforming baselines such as GR00T N1.5, $\pi_{0}$, $\pi_{0}$-FAST, and Magma. With fine-tuning on the same RT-1 subset of OXE, MolmoAct-7B-D improved to 71.6%, exceeding Magma by 3.2% as shown in Table 1. These results indicate that MolmoAct is both an effective zero-shot generalist and a strong initialization for fine-tuned deployment.
@@ 5.2 Fast Adaptation of MolmoAct in Post-training
Evaluation Setups and Baselines. We evaluate MolmoAct in both simulation and real-world settings to assess its fast adaptation after post-training. In simulation, we evaluate on the LIBERO simulation benchmark (Liu et al. (2023a)), which consists of a Franka Emika Panda arm in simulation with demonstrations containing front and wrist view camera images (256256 px), language instructions, and delta end-effector pose actions. We follow prior works (Kim et al. (2024)) and evaluate on four task suites – LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long – each with 500 demonstrations across 10 tasks. Following (Kim et al. (2024)), we trained on a modified dataset that filtered out no-op actions and unsuccessful demonstrations. Moreover, we set action chunk size to $K=8$ for evaluation on each task suite and execute full chunks before redoing action reasoning. We fine-tune MolmoAct-7B-D using Low-Rank Adaptation (LoRA) and compared to state-of-the-art generalist autoregressive policies, such as TraceVLA (Zheng et al., 2024), OpenVLA (Kim et al., 2024), SpatialVLA (Qu et al., 2025), $\pi_{0}$-FAST (Black et al., ), CoT-VLA (Zhao et al., 2025), WorldVLA (Cen et al., 2025), ThinkAct (Huang et al., 2025), and NORA-AC (Hung et al., 2025).